# PAC in DESI. I. Galaxy Stellar Mass Function into the $10^{6}{\rm M}_{\odot}$ Frontier

### v3

@@ Acknowledgements
K.X. is supported by the funding from the Center for Particle Cosmology at U Penn. Y.P.J. is supported by NSFC (12133006, 11890691, 11621303), by grant No. CMS-CSST-2021-A03, and by 111 project No. B20019. Y.P.J. gratefully acknowledges the support of the Key Laboratory for Particle Physics, Astrophysics and Cosmology, Ministry of Education. S.B. is supported by the UK Research and Innovation (UKRI) Future Leaders Fellowship [grant number MR/V023381/1]. W.W. is supported by NSFC (12273021) and the National Key R&D Program of China (2023YFA1605600, 2023YFA1605601). This work made use of the Gravity Supercomputer at the Department of Astronomy, Shanghai Jiao Tong University. This work used the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/K00042X/1, ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the National e-Infrastructure.

### v4

@@ Acknowledgements
K.X. is supported by the funding from the Center for Particle Cosmology at U Penn. Y.P.J. is supported by NSFC (12133006, 11890691, 11621303), by grant No. CMS-CSST-2021-A03, and by 111 project No. B20019. Y.P.J. gratefully acknowledges the support of the Key Laboratory for Particle Physics, Astrophysics and Cosmology, Ministry of Education. S.B. is supported by the UK Research and Innovation (UKRI) Future Leaders Fellowship [grant number MR/V023381/1]. $W.$ is supported by NSFC (12273021) and the National Key R&D Program of China (2023YFA1605600, 2023YFA1605601). This work made use of the Gravity Supercomputer at the Department of Astronomy, Shanghai Jiao Tong University. This work used the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/K00042X/1, ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the National e-Infrastructure.